Find the vapor pressure of a solution containing 35.0 g calcium chloride (CaCl2) in 750.0 g water. Note that P^0 water = 47.1 torr at 37.0 degrees C.

1
Expert's answer
2022-03-08T03:10:04-0500

Solution:

Moles CaCl2: 35.0 g / 110.98 g/mol = 0.315 mol

Moles H2O: 750.0 g / 18.02 g/mol = 41.62 mol

Mole fraction H2O: 41.62 / (3 x 0.315 + 41.62) = 0.9778

CaCl2 = Ca2+ + 2Cl- (i = 3, 3 ions)

The vapor pressure of the solution at 370C:

0.9778 x 47.1 = 46.05 torr